OpenStudy (anonymous):

giải pt sin3x / cos3x -1 =0

OpenStudy (anonymous):

What we are to do here ?? Finding the solution for x??

OpenStudy (anonymous):

What does that pt stand for ????

myininaya (myininaya):

That is \[\frac{\sin(3x)}{\cos(3x)}-1=0\] , right?

OpenStudy (anonymous):

Yes I think..

myininaya (myininaya):

I would add 1 to both sides and recall that sin(u)/cos(u)=tan(u) Where u=??? I will let you reply back before I add anything else @nhỏ